INAF queue observing, 2022 Apr 23 - Apr 24

Observer: A. Cardwell
Telescope Operator: S. Allanson
Astronomers from Italy: Felice Cusano, Ester Marini
Instrument: LBC

Summary

Due to the Covid-19 pandemic situation these observations have been performed by LBTO personel
through the indications of LBT Italian observers (F. Cusano, A. Rossi, S. Paiano and E. Marini) connected from Italy.

At sunset the sky is clear. We do not have internet connection to the mountain as the high wind of yesterday damaged the microwave antenna.
For this reason we can not do LUCI-AO observations as there is no AO support. The support astronomer A. Cardwell is driving to the mountain to perform regular observations with LBC. We start with LBC-bino.

Night duration (12 degree - twilight) 02:55 - 11:55 UT = 9 h 00 m
Observing time: 9h 00min (100%)*
Weather time loss: 0h 00min (0%)*
Technical time loss: 0h 00min (0%)*

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibration flat fields LBC-Bino       r+r
Palazzi SLSN2020abjc LBC-Bino ug+ri     completed
Nardiello NGC5024 LBC-Bino U-Bessel + V-Bessel 1.68h+1.68h   UV-3,4,5 done, fwhm~1-1.2",~0.8-1" (UV-5), for UV-5 see notes b
Nardiello NGC5024 LBC-Bino B-Bessel + I-Bessel 0.5h+0.5h   BI-4,5 done, seeing~0.8"(BI-5),~1"(BI_4), for BI-5 see note a
vito J1641 LBC-Bino g+i sloan 0.5h+0.5h   seeing~0.8-1", 1.2" from 2/3 part of the observation to the last two exposures when it was ~1"
Nardiello NGC6341 LBC-Bino B-Bessel + I-Bessel 0.5h+0.5h   BI-1,2 done but in BI-1 seeing very unstable, from 1.1" to 1.5", ~0.8" in BI-2
Giunta AO_I LBC-Bino g+i sloan 0.12h+0.12h   Slight jump in initial exposure, we repeated it
Giunta AO_J LBC-Bino g+i sloan 0.09h+0.09h   excellent seeing
Nardiello NGC6341 LBC-Bino U-Bessel + V-Bessel 1.12h+1.12h   UV-1,2 done, seeing ~1" in UV-1 (satellite trail in lbcr.095806), ~0.8" in UV-2
Giunta twilight field LBC-Bino g+i,r+z 0.13h+0.13h   we run the scrip twice and stopped at 50% of the second (20k background)
calibration flat fields LBC-Bino       r+z PA=0 PA=180
Detailed log by Andrew Cardwell at this link:

https://drive.google.com/file/d/15GaSpu_ysCK2qjH-YKlDMhLlsLMe6GTJ/view?usp=sharing

Notes

a- Slight image jump in lbcb.20220424.060013 and lbcr.20220424.060011. Possibly due to TMS?

b- lbcb.063505 is trailed, Andrew and Steve believe it was a guiding (or more likely lack of guiding) issue, see the log for details.
Topic revision: r7 - 26 Apr 2022, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback